OpenStudy (zzr0ck3r):

Find a power series representation of the function \(f(x) =\large \frac{x}{2x^2+1}\)

OpenStudy (zzr0ck3r):

@satellite73 @ganeshie8 @mathstudent55

OpenStudy (zzr0ck3r):

they are doing these things where they are putting these in the form \(\frac{a}{1-r}\)m and then putting them into geometric sums. But I cant figure out how to do it with this one.

OpenStudy (aum):

For |x| < 1, \[\large \frac{1}{1-x} = \sum_{n=0}^{\infty} x^n,~~~~~~|x| \lt 1 \\ \large \frac{x}{2x^2+1} = \frac{x}{1-(-2x^2)} = x * \sum_{n=0}^{\infty} (-2x^2)^n = \sum_{n=0}^{\infty} (-2)^nx^{2n+1} \]

OpenStudy (aum):

The second line above is the standard sum of an infinite geometric series: 1 + x + x^2 + x^3 + .... = 1 / (1-x) for -1 < x < 1 For the given problem, the interval of convergence is: \( \large |-2x^2| < 1 \\ \large -\frac{\sqrt{2}}{2} \lt x \lt \frac{\sqrt{2}}{2} \)

OpenStudy (anonymous):

asdf_movie

OpenStudy (anonymous):

hi girl

OpenStudy (anonymous):

hi samsonite

OpenStudy (anonymous):

what u doin

OpenStudy (anonymous):

i knew that

OpenStudy (anonymous):

what ur number

OpenStudy (anonymous):

naw jk can u help me

OpenStudy (anonymous):

math pre al

OpenStudy (anonymous):

Which graph shows the solution for the inequality